part 1: A window has a glass surface of 3810 cm2 and a thickness of 8.8...

part 1:

A window has a glass surface of 3810 cm2 and a thickness of 8.8 mm. Find the rate of energy transfer by conduction through this pane when the temperature of the inside surface of the glass is 70◦F and the outside temperature is 80 ◦F. Assume the thermal conductivity of window glass is 0.8 J/s · m · ◦ C . Answer in units of kW.

Part 2:

Find the rate of energy transfer for the same inside temperature and an outside temperature of 0◦F. Answer in units of kW.
Answer #1

Part 1:

we use Fourier law of heat conduction,

Q=kA \frac{\Delta T}{\Delta x}

Tin= 70 F=> 21.11 degree C,

Tout= 80 ◦F=> 26.6 degree C,

A= 3810 cm2 => 0.381 m2

k= 0.8 J/s · m · ◦ C.

Q= 0.8 * 0.381*(26.6-21.11)/8.8*10-3

Q= 190 J/sec,

Q= 0.190 KW.

Part 2

for 0 F, the temperature in celcius is, -17.7 degree C.

Q= 0.8 * 0.381*(21.11-(-17.7))/8.8*10-3

Q= 1344.23 J/sec,

Q= 1.34 KW.